OpenStudy (anonymous):

Anyone need help come ask me ill even medal u 4 ur question

OpenStudy (kenialove):

@Peaches__15 can u help me ?

OpenStudy (anonymous):

with wat

OpenStudy (kenialove):

I posted a question that I need help with

OpenStudy (amtran_bus):

Peaches!

OpenStudy (anonymous):

Complete the square in the quadratic function f given by f(x) = 2x2 - 6x + 4 Can you help with this

OpenStudy (anonymous):

kk hold on 1 at a time lol

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

(x) = 2x² – 6x + 4 Is there anything we can pull out of all three terms? I believe we can pull out a 2. f(x) = 2(x² – 3x + 2) Now factor, if possible: f(x) = 2(x – 2)(x – 1) Now, set the function equal to 0. 0 = 2(x – 2)(x – 1) Divide both sides by 2 to get rid of that 2: 0 = (x – 2)(x – 1) Now, set each part equal to 0: x – 2 = 0 x = 2 x – 1 = 0 x = 1 Therefore, you have two solutions: x = 1, 2 Hope I helped! plzz medal me if I help u lol

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

@AmTran_Bus wat do u need help with